### v1

# ELUTQ: Efficient LUT-Aware Quantization for Deploying Large Language Models on Edge Devices

Abstract
The deployment of Large Language Models on CPU-based edge devices is crucial for enabling on-device intelligence and broadening the reach of AI applications. However, such deployment remains challenging due to the limited memory and computational resources typical of these devices. When performing inference on edge CPUs, memory usage and latency are two primary bottlenecks. Weight quantization effectively reduces memory consumption, yet existing hardware-friendly methods often rely on uniform quantization, which suffers from poor weight-distribution fitting and high dequantization overhead under low-bit settings. To address these issues, we propose ELUTQ, an efficient quantization framework featuring a novel quantization format termed Hierarchical Linear Quantization (HLQ). HLQ is designed to better capture the statistical characteristics of weights without increasing the computational cost of Bit-serial LUT-based GEMM operations, thereby eliminating dequantization overhead. As a fundamental quantization scheme, HLQ is orthogonal to existing quantization algorithms and can be seamlessly integrated into various quantization pipelines. To enable efficient deployment on edge devices, ELUTQ designs high-performance CPU kernels to support end-to-end inference. Extensive experiments demonstrate the effectiveness of our approach. For the LLaMA3-8B model, when combined with the post-training quantization framework, HLQ enhances uniform quantization by achieving approximately 8% perplexity reduction at 3-bit precision and 85% perplexity reduction at 2-bit precision, with quantization completed within one hour. When combined with efficient finetuning techniques, HLQ further improves perplexity under the 2-bit configuration and completes quantization in about two hours. In terms of inference efficiency, under a 4-thread configuration with batch size = 1, our 2-bit quantized LLaMA2-7B model achieves a throughput of over 25 tokens per second on an Apple M2 chip. All the code is available at https://github.com/Nkniexin/ELUTQ.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across diverse tasks, including natural language understanding, image recognition, and multimodal reasoning. Traditionally deployed in cloud environments with abundant computational resources, these models are now increasingly being adapted for edge devices, such as smartphones, IoT systems, and autonomous vehicles, to meet growing demands for low-latency inference, privacy preservation, and real-time intelligent services.
Unlike high-performance GPU servers, on-device hardware typically operates under stringent resource constraints, characterized by limited memory capacity and computational power. These devices predominantly employ ARM or x86 CPUs, which offer restricted vectorization support and limited parallelism. To address these challenges, model quantization (Frantar et al., 2022; Xiao et al., 2023; Lin et al., 2024; Kim et al., 2023; Shang et al., 2023; Chen et al., 2024b) has emerged as a widely adopted compression technique, where high-precision weights are mapped to discrete integer values and stored using low-bit representations. This approach drastically reduces memory footprint while preserving model accuracy. Recent advances demonstrate that 8-bit weight quantization achieves near-lossless performance (Xiao et al., 2023; Yao et al., 2022). Furthermore, 4-bit quantization techniques (Frantar et al., 2022; Lin et al., 2024; Shao et al., 2023; Kim et al., 2023; Chen et al., 2024c) typically incur less than 3% accuracy degradation, with ongoing research further improving robustness. However, activation quantization remains more challenging due to the prevalence of outlier values. While some studies (Xiao et al., 2023; Liu et al., 2024) explore low-bit activation quantization, most practical implementations retain activations in FP16 precision (e.g., W4A16, W3A16) to ensure accuracy. Meanwhile, some research (Chen et al., 2024a; Shang et al., 2023; Huang et al., 2024) focuses on extreme low-bit weight quantization, pushing the boundaries of efficiency without sacrificing model quality.
Figure 1: Overview of ELUTQ.
Quantization can be categorized as uniform (Frantar et al., 2022; Lin et al., 2024; Shao et al., 2023; Chen et al., 2024c) or non-uniform (Chee et al., 2023; Park et al., 2024; Kim et al., 2023; Xu et al., 2023; Zhao & Yuan, 2025) depending on whether the quantization intervals are equal. Uniform quantization divides the weight space into equal intervals, making it hardware-friendly and efficient for acceleration (Lin et al., 2024; Frantar et al., 2022). However, as noted in (Dettmers et al., 2023a; Kim et al., 2023), this approach poorly matches the bell-shaped distribution of typical weights, incurring substantial approximation error. Non-uniform quantization addresses this by adaptively allocating bins, via clustering (Kim et al., 2023; Zhao & Yuan, 2025) or codebooks (Chee et al., 2023) to better fit the weight distribution. While this improves representational efficiency, it often sacrifices hardware compatibility due to irregular memory access patterns. To address these limitations, We propose Hierarchical Linear Quantization, a non-uniform quantization format that reduces weight quantization error while maintaining hardware compatibility.
Many large model quantization methods aim to efficiently complete model quantization, where efficiency primarily refers to time and memory consumption, particularly video memory (VRAM), which is typically more constrained and expensive than CPU memory. Achieving quantization with low memory usage and short processing time is critical, as it enables running quantization algorithms even on consumer-grade GPUs such as the RTX 3090 or RTX 4090. Most post-training quantization approaches (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023; Huang et al., 2024; Shang et al., 2023; Zhao & Yuan, 2025; Kim et al., 2023) can complete quantization quickly and with modest memory requirements. These train-free methods operate without retraining and typically quantize the model layer by layer. For example, GPTQ quantizes one linear layer at a time. With sufficient CPU memory, even large models such as LLaMA2-70B can be quantized on a single A6000 GPU. In contrast, efficient finetuning–based methods (Shao et al., 2023; Dettmers et al., 2023b, a; Chee et al., 2023; Li et al., 2023; Xu et al., 2023; Chen et al., 2024b) require partial retraining, which increases time cost but often mitigates VRAM pressure by freezing most parameters, training only a small subset, or performing quantization in batches. This strategy allows quantization to be completed within limited video memory. In this paper, we combine our proposed Hierarchical Linear Quantization with efficient quantization techniques to enhance model accuracy while maintaining low VRAM consumption and short quantization time, thereby further advancing on-device intelligence.
Although weight quantization reduces memory footprint, traditional methods require dequantization to higher precision (8-bit/FP16) for computation, introducing significant overhead that can paradoxically slow down inference at low bit-widths. Recent advances (Wei et al., 2025; Park et al., 2025, 2022) address this by replacing standard GEMM with lookup table (LUT)-based operations that implement generalized FP-INT multiplication, eliminating dequantization while achieving both linear latency reduction with bit-width and improved energy efficiency. FLGLUT (Park et al., 2025) accelerates these operations on GPUs and T-MAC (Wei et al., 2025) optimizes for CPUs via SIMD instructions. We enhance this paradigm through a pure C++ kernel redesign that specifically supports our novel Hierarchical Linear Quantization format, maintaining cross-platform compatibility while optimizing for our method’s unique requirements.
Figure  1 shows the ELUTQ design. Our contributions are as follows.
  • We propose Hierarchical Linear Quantization (HLQ), a novel non-uniform quantization format that provides greater flexibility in weight representation compared to uniform quantization, while enabling efficient matrix computation through LUT-based GEMM.
  • We integrate HLQ into existing efficient quantization methods, including post-training quantization and efficient fine-tuning techniques, demonstrating that HLQ is orthogonal to most existing methods and can significantly enhance model performance under low-bit settings, without introducing noticeable memory or time overhead to the quantization pipeline.
  • We design an efficient CPU kernel tailored for the HLQ format, which enables high-performance matrix operations on edge devices.
  • We introduce ELUTQ, a unified quantization framework built upon HLQ. ELUTQ incorporates common quantization optimization techniques and includes a fully C++-implemented inference runtime, which supports end-to-end deployment of quantized models on edge devices.
2 Related works
2.1 Model Quantization
2.1.1 Uniform Quantization.
The uniform quantization formula is as follows:
Quantization:
${\mathrm{W}_{int}=\mathrm{clamp}}(\lfloor\frac{\mathrm{W}}{s}\rceil+z,0,2^{q}-1).$ (1)
Here, $\lfloor\cdot\rceil$ denotes the rounding operation, $q$ is the quantization bit width, $s$ is the scale factor, $z$ is the zero-point, $\mathrm{W}$ represents the original weights, and $\mathrm{W}_{int}$ denotes the quantized integer weights.
Dequantization:
$\mathrm{\hat{W}}=(\mathrm{W}_{int}\ -\ z)\cdot s,$ (2)
where $\mathrm{\hat{W}}$ represents the dequantized weights. Uniform quantization is hardware-friendly, as dequantization can be implemented with simple multiplication. However, its representational capability is limited because it maps weights into a uniformly spaced range. Moreover, in modern computing systems, the smallest storage and computation unit is typically 8 bits. Therefore, for low-bit settings (e.g., 2-bit or 3-bit), the quantized integer weights $\mathrm{\hat{W}}$ must first be dequantized into 8-bit or 16-bit formats before computation, which often introduces non-negligible computational overhead.
2.1.2 Efficient Quantization.
Efficient quantization aims to complete model quantization with minimal time and memory consumption. It is generally categorized into two types: post-training quantization (PTQ) and efficient finetuning. PTQ requires only a small calibration dataset and can be completed within a relatively short time. Several research streams have advanced PTQ for Large Language Models. Some works, such as GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024), focus on weight-only quantization, demonstrating minimal performance degradation (e.g., less than 3% perplexity increase) with 4-bit uniform quantization. Other approaches (Dettmers et al., 2023b; Kim et al., 2023) differentiate weights by their significance, storing a subset of important weights in higher precision. Furthermore, studies including SqueezeLLM (Kim et al., 2023) and GANQ (Zhao & Yuan, 2025) observe that weights in LLMs often follow a bell-shaped distribution and subsequently propose non-uniform quantization methods based on k-means clustering. Beyond weight quantization, activation quantization has also been explored. Methods like SmoothQuant (Xiao et al., 2023) and LLM.int8() (Dettmers et al., 2022) have achieved notable results in the challenging W8A8 weight-activation joint quantization setting. Compared with post-training quantization, efficient finetuning typically requires a little more data and time to search and adjust model parameters. OmniQuant (Shao et al., 2023) searches quantization parameters block by block and is the first to achieve promising results under 2-bit settings. PB-LLM (Shang et al., 2023) employs a feature segmentation strategy to achieve competitive performance under 2-bit weight quantization. Meanwhile, DB-LLM (Chen et al., 2024b) decomposes 2-bit quantized weights into two independent sets of binary weights and further utilizes distillation to enhance model performance, albeit at the cost of introducing substantial fine-tuning overhead. BiLLM (Huang et al., 2024) pushes the boundary further by leveraging weight distribution characteristics to achieve an average bit-width of approximately 1.11 bits.
Figure 2: Illustration of three computation paradigms for weight-only quantized matrix multiplication.
Figure 3: Detail pipeline of Bit-serial Lut-based GEMM.
2.2 LUT-Based GEMM
There are two primary computational paradigms for LUT-based GEMM, as illustrated in the Figure 2.
Figure 2(b) depicts the first implementation scheme of LUT-based GEMM. Its core idea is to directly store the high-precision weight values corresponding to low-bit integer weights in the LUT. Consequently, during dequantization, high-precision weights can be reconstructed via direct LUT accesses instead of computationally expensive arithmetic operations, significantly reducing the overhead of dequantization. It should be noted that, in this paradigm, although weights are reconstructed into a high-precision format via the LUT, they maintain the same numerical precision (e.g., both FP16) as the activations during the matrix multiplication.
Figure 2(c) illustrates another GEMM paradigm termed Bit-serial LUT-based GEMM. Unlike the scheme in Figure 2(b), this method utilizes the LUT to store all possible dot products between an activation vector and single-bit weights. The detail computational procedure is shown in Figure 3. First, a $q$-bit quantized weight matrix $\mathrm{W}_{int}$ is decomposed into $q$ single-bit matrices $\{W_{0},W_{1},...,W_{q-1}\}$ offline, where each element is either 0 or 1, representing the respective bit planes of the original weights. For example, for integer values (9, 7, 6, 3) with binary representations (1001, 0111, 0110, 0011), the matrix for the lowest bit is (1, 1, 0, 1), and the matrix for the highest bit is (1, 0, 0, 0). This decomposition is performed offline, incurring no runtime overhead. During inference, for an activation vector of the group size $g$, the system precomputes the dot products between this activation vector and all $2^{g}$ possible combinations of single-bit weights, storing the results in the LUT. Thus, the original matrix computation requiring high-precision multiply-accumulate operations is simplified into highly efficient table lookups followed by summation. This paradigm has been demonstrated to offer high computational efficiency and energy efficiency (Park et al., 2025; Wei et al., 2025). For example, FIGLUT (Park et al., 2025) optimized the table structure for GPU architectures to avoid bank conflicts, while T-MAC leveraged CPU vectorized lookup instructions (AVX2/NEON) to enable efficient LUT operations on CPUs.
The HLQ method is built upon the Bit-serial LUT-based GEMM computational paradigm. In contrast to prior works like FIGLUT and T-MAC, which primarily focused on designing efficient computational kernels for this paradigm, our work emphasizes optimization at the quantization algorithm level. The objective is to enhance the accuracy of low-bit quantized models, thereby achieving a synergistic improvement in both accuracy and efficiency within this highly efficient computational paradigm.
2.3 Inference Framework
Many frameworks aim to enable efficient inference of quantized models across a wide range of hardware platforms. For GPUs, vLLM (Kwon et al., 2023) employs the PageAttention technique to optimize key-value (KV) memory management and supports quantization formats such as GPTQ and AWQ. SGlang (Zheng et al., 2024) reduces response latency through shared prefix requests and efficient caching strategies. In addition, frameworks like TensorRT-LLM (NVIDIA, 2023) and MLC-LLM (Chen, 2023) have also been developed for GPU-optimized inference. For CPUs and other edge processors, llama.cpp (Gerganov, 2023) is a lightweight framework implemented entirely in C++. Although its inference speed is slower compared to GPU-accelerated solutions and thus not suitable for large-scale online services, it is well-suited for edge computing, IoT, and low-throughput scenarios, providing a practical solution for basic inference in GPU-free environments.
3 Motivation
3.1 Bell-Shaped Distribution of Weights
As noted in prior studies (Kim et al., 2023; Dettmers et al., 2023a; Zhao & Yuan, 2025), weights often exhibit a bell-shaped distribution. Uniform quantization, which maps weights to a uniformly spaced grid, struggles to accurately capture such distributions, especially under low-bit settings. Figure 4 shows that the weight distribution of an out_channel in the up_projection layer of the LLaMA3-8B model closely follows a Gaussian pattern, with values heavily concentrated near zero. This characteristic, commonly observed across various models, highlights the inherent limitation of uniform quantization in effectively representing natural weight distributions. Consequently, non-uniform quantization strategies are often necessary to achieve higher fidelity in low-bit compression.
Figure 4: (Top) The weight distribution of one output channel in a up_proj of LLaMA3-8B. (Bottom) MSE loss for weight comparsion between Uniform and HLQ.
3.2 Optimization for Bit-Serial LUT-Based GEMM
For ordinary LUT-based GEMM algorithms shown in figure 2(b), previous efforts such as SqueezeLLM (Kim et al., 2023) and GANQ (Zhao & Yuan, 2025) have achieved significant improvements in algorithmic performance for 3-bit and 4-bit quantization, along with the development of highly efficient CUDA kernels for accelerated inference. However, these implementations still lack support for CPU devices. Regarding Bit-serial LUT-Based GEMM, existing works including T-MAC (Wei et al., 2025), LUT-GEMM (Park et al., 2022), and FIGULT (Park et al., 2025) have primarily focused on optimizing kernel design to support efficient uniform quantization schemes, yet they do not address algorithmic enhancements for improving quantization accuracy. This limitation restricts the broader application of such kernels. To bridge this gap, our work proposes using HLQ to boost quantization accuracy at the algorithm level, while leveraging Bit-Serial LUT-Based GEMM to enable efficient inference on CPU devices.
3.3 Why Choose Bit-Serial LUT-Based GEMM for Edge Devices
Edge devices are predominantly based on CPU architectures, which offer significantly lower programmibility compared to GPU architectures. This limitation prevents developers from customizing lookup tables to implement LUT-based GEMM shown in figure 2(b). However, CPUs with x86_64 or ARM architectures provide a set of vectorized table-lookup instructions, such as _mm256_shuffle_epi8 and vqtbl1q_u8, which naturally align with the computational paradigm of Bit-serial LUT-based GEMM. Prior work such as T-MAC (Wei et al., 2025) has demonstrated that these instructions can be leveraged to enable efficient execution of Bit-serial LUT-based GEMM on edge devices. This observation motivates our work to bridge the algorithm-kernel co-design gap by combining high-accuracy non-uniform quantization at the algorithm level with a Bit-serial LUT-based inference framework tailored for common CPU instruction sets.
4 Methodology
In this section, we first present the proposed Hierarchical Linear Quantization method. Then, we discuss its integration with existing efficient quantization techniques. Finally, we describe the memory organization strategy and LUT design that enable efficient inference on edge devices.
4.1 Hierarchical Linear Quantization
As discussed previously, most PTQ methods, such as GPTQ and AWQ, adopt uniform quantization to facilitate hardware acceleration. However, uniform quantization cannot effectively represent the distribution of weights. Motivated by the empirical observation that weight distributions in LLMs tend to follow a bell-shaped curve (Kim et al., 2023), inspired by Binary Coding (Xu et al., 2018), we propose Hierarchical Linear Quantization as an alternative approach. Specifically, for an $n$-dimensional weight vector $\mathrm{W}$, its $q$-bit quantized representation is denoted as $\mathrm{\hat{W}}$:
$\mathrm{\hat{W}}=\sum_{j=0}^{q-1}s_{j}\cdot b_{j}+z.$ (3)
Here, $b_{j}$ is a binary vector $\in\{0,1\}^{n}$$s_{j}\in R$ is the quantization scale, and $z\in R$ is the zero-point. Similar to uniform quantization, given $s$ and $z$, we present the quantization process of HLQ. For a $q$-bit quantization, we first generate all possible binary combinations, which form a codebook denoted by $C\in\{0,1\}^{2^{q}\times q}$. Based on this codebook, we construct a candidate set:
$V=s\times C^{T}+z$ (4)
Then, we define Quantization and Dequantization as follows:
Quantization:
[…]
Dequantization:
$\mathrm{\hat{W}}=s\times B^{T}+z$ (6)
It is worth noting that HLQ neither requires weight reconstruction nor introduces any additional factors, making it highly generalizable. It can be seamlessly integrated with various quantization methods. As a form of non-uniform quantization, it not only offers significant advantages in reducing weight representation error (see Section 4.2), but also introduces no additional computational overhead to Bit-serial LUT-based GEMM (see Section 9).
4.2 Weight Error Reduction via Hierarchical Linear Quantization
In this subsection, we introduce how to use Hierarchical Linear Quantization to reduce weight quantization error. Let $\mathrm{\hat{W}}=\mathrm{HLQ}(\mathrm{W};s,z)$, The objective is to find the optimal set of $q$ scales and a zero-point $z$ to represent $\mathrm{W}$. The optimization objective can be formulated as:
[…]
For this optimization problem, there are two possible approaches: a heuristic alternating optimization method and a gradient-based search method. Both approaches have their own advantages and limitations. The alternating optimization method features a simple computation flow and can efficiently obtain a locally convergent solution. In contrast, the gradient-based approach relies on backpropagation, which often achieves better global optimization results but is computationally more expensive. Here, we just introduce Gradient-based search method. Alternating optimization can be seen in Appendix A
The gradient-based search updates the quantization parameters by directly computing the MSE and propagating the gradients. At initialization, we set $z=\min(\mathrm{W})$, and the initial value of s is set to the scale factor used in uniform quantization. Specifically, let $\Delta=\frac{\max(\mathrm{W})-\min(\mathrm{W})}{2^{q}-1}$, For a $q$-bit quantization, the initial value of s is set to $s=[\Delta,2\Delta,...2^{q-1}\Delta]$. The workflow of the search process is illustrated in Algorithm 1. Here, we adopt an approximately differentiable optimization strategy to jointly update the scales and zero-points. The core idea is to decouple the quantization process into two stages: discrete selection and continuous reconstruction, thereby circumventing the gradient issues inherent in discrete operations.
  • Forward Pass: the optimal combination of discrete codewords is determined via nearest-neighbor search. Subsequently, this combination is used together with learnable scaling factors and zero-points to reconstruct the quantized weights through a continuous weighted summation.
  • Backward Pass: During backpropagation, gradients cannot flow through the argmin-based discrete selection step. However, they can naturally propagate back through the continuous reconstruction step to update the scaling factors and zero-points. This enables the continuous parameters to be effectively optimized via standard gradient descent.
To ensure training stability, we enforce a non-negativity constraint on the scaling factors and apply dynamic range clipping to the zero-point values, which effectively prevents training divergence. From an optimization perspective, this approach is conceptually aligned with the straight-through estimator (STE) (Liu et al., 2022) philosophy. However, a key distinction lies in the fact that instead of relying on handcrafted gradient approximation rules, we carefully design the quantization expression itself to allow gradients to flow naturally to the continuous parameters, resulting in a more elegant approximation of gradient flow.
Figure 4 shows that under these two optimization methods, HLQ significantly reduces the quantization error compared to uniform quantization, with the MSE of uniform quantization at the 1e-4 level, while HLQ decreases it to the 1e-6 level.
Algorithm 1 Gradient-Based Optimization for Hierarchical Linear Quantization
0: Weight matrix $\mathrm{W}\in R^{n\times k}$, bit width q, Group size g, Max iterations $T_{max}$, Tolerance $\varepsilon$
0: scales $s\in R^{n\times\frac{k}{g}\times q}$, zero points $z\in R^{n\times\frac{k}{g}}$
1: Reshape W into groups $\mathrm{\widetilde{W}}\in R^{n\times\frac{k}{g}\times g}$
2: Calculate uniform quantization scale : $\Delta\leftarrow\frac{\max(\mathrm{\widetilde{W}})-\min(\mathrm{\widetilde{W}})}{2^{q}-1}$
[…]
4: Generate binary combinations $C$
5:for $t\leftarrow 1$ to $T_{max}$ do
6:  $\mathrm{\hat{W}^{(t)}}\leftarrow\mathrm{\widetilde{W}}-z^{(t-1)}$
7:  $V^{(t)}\leftarrow s^{(t-1)}\times C^{T}$ # Matrix product
8:  $k^{*}\leftarrow\arg\min_{k}\|\mathrm{\hat{W}^{(t)}}-V^{(t)}_{k}\|$
9:  $B^{(t)}\leftarrow C_{k^{*}}$ # Choose best binary combination
10:  $\mathrm{\hat{W}^{(t)}}\leftarrow s^{(t-1)}\times{B^{(t)}}^{T}$
11:  $\mathrm{\hat{W}^{(t)}}\leftarrow\mathrm{\hat{W}^{(t)}}+z^{(t-1)}$
12:  $L\leftarrow mean((\mathrm{\widetilde{W}}-\mathrm{\hat{W}^{(t)}})^{2})$
13:  $s^{(t)},\ z^{(t)}\leftarrow Backpropagate(L)$
14:  if $L-preL<\varepsilon$ then
15:   break
16:  end if
17:  $perL\leftarrow L$
18:end for
4.3 Post-Training Quantization for HLQ
Here, using GPTQ as an example, we illustrate how HLQ can be employed to enhance existing post-training quantization methods. GPTQ, a widely adopted PTQ approach, offers strong hardware efficiency due to its use of group-wise uniform quantization and can quantize models within a very short time. However, it suffers from significant accuracy degradation under 2-bit quantization. By integrating HLQ into the GPTQ quantization pipeline, we effectively improve its performance in both 2-bit and 3-bit settings.
Figure 5 illustrates the integration of HLQ into the GPTQ pipeline. As shown in Figure 5(a), the standard GPTQ process partitions the weights into multiple column blocks and recursively quantizes each block column by column, using unquantized columns to compensate for quantization errors. After a block is fully quantized, subsequent unquantized blocks are leveraged to further correct accumulated errors. In contrast, Figure 5(b) presents our HLQ-GPTQ procedure. Instead of recursive column-wise quantization, our method directly applies HLQ to each column block as a whole, while retaining the error compensation mechanism through subsequent blocks. In this PTQ pipeline, we only replace the block-wise quantization step with HLQ, keeping all other components identical to GPTQ. Despite this straightforward modification, it substantially enhances GPTQ’s performance in low-bit quantization scenarios (see Section 5.1.4).
It should be noted that GPTQ commonly recommends reordering weight columns based on the diagonal elements of the Hessian matrix to achieve improved quantization performance. While our approach is theoretically compatible with this reordering strategy, it incurs additional activation reordering overhead during inference. Furthermore, prior research (Frantar & Alistarh, 2023) has indicated that such Hessian-based reordering may cause overfitting, particularly under low-bit quantization settings. Therefore, we omit this reordering optimization in our final design.
Although integrating HLQ introduces additional parameter search overhead to GPTQ, this cost remains within a controllable range. Specifically, HLQ-GPTQ can still quantize the Llama2-7B model within half an hour using only 8 GB of GPU memory, thereby retaining the time and memory efficiency characteristic of PTQ methods (see Section 5.3). It is worth noting that under 2-bit quantization, the performance of HLQ-GPTQ still lags behind state-of-the-art PTQ methods such as DB-LLM (Chen et al., 2024b). However, this performance gap primarily arises from fundamental differences in methodological complexity. Methods like DB-LLM typically incorporate additional procedures such as knowledge distillation, which require significantly more computational and memory resources and often take tens of times longer to complete quantization. In contrast, HLQ-GPTQ preserves the simplicity and hardware friendliness of the PTQ paradigm. More importantly, HLQ serves as a fundamental quantization format that is largely orthogonal to most post-training optimization techniques. This implies that by integrating existing PTQ refinements, the performance of HLQ under low-bit settings can be further improved, leaving ample room for future research.
Figure 5: Standard GPTQ VS HLQ-GPTQ.
4.4 Efficient Finetuning for HLQ
Previous quantization methods (Shao et al., 2023; Chen et al., 2024b; Li et al., 2021; Chen et al., 2024c) typically improve quantization performance by performing block-wise search for optimal quantization parameters or end-to-end finetuning. However, most of these methods are designed for uniform quantization and are not directly applicable to HLQ. As shown in figure  6, We propose an efficient finetuning scheme tailored for HLQ, which consists of two parts: block-wise reconstruction and end-to-end tuning.
The block-wise reconstruction aims to minimize the block output error. We adopt the same objective but replace uniform quantization with HLQ :
[…]
where $\mathcal{F}$ denotes the mapping function of a Transformer block, $\mathbf{W}$ and $\mathbf{X}$ are the full-precision weights and activations, and $\mathrm{HLQ}$ represents the Hierarchical Linear Quantization function. To jointly consider local and block-level errors, we divide the block-wise reconstruction into two stages. In the first stage, we search for the optimal $\mathbf{W}_{int},s,z$ for each linear layer within the block using the technique described in Algorithm 1. In the second stage, we fix $\mathbf{W}_{int}$ and only optimize the scale $s$ and zero-point $z$. For each linear layer, this two-strategy preserves some local information while leveraging block-level context to further refine the parameters. Additionally, this two-stage scheme is also computationally efficient. In the first stage, parameter initialization for each linear layer is independent, allowing for batch-wise optimization. In the second stage, with $\mathbf{W}_{int}$ fixed, there is no need to access the original full-precision weights, further reducing computation and memory footprint. The experiments in Section 5.3 demonstrate the efficiency of this scheme.
After completing the block-wise reconstruction, we introduce end-to-end tuning, a training-based refinement approach. In this stage, we perform end-to-end optimization on the model using a calibration dataset. This technique is also adopted in EfficientQAT (Chen et al., 2024c), and following their practice, we update only the quantization scale $s$ during training. To avoid overfitting, we limit the tuning process to only 1–2 epochs.
Figure 6: The pipeline of efficient finetuning for HLQ.
Figure 7: Rearrange weight for ARM NEON’s 128-bit registers to ensure memory access continuity and improve decoding speed during runtime.
4.5 Deploy with Edge Framework
In this subsection, we present the core design of the C++ inference framework tailored for Hierarchical Linear Quantization.
4.5.1 Weight-Rearrange.
Conventional dequantization-based inference frameworks typically store weights in either row-major or column-major order. Some works, such as AWQ, employ interleaved weight storage for 4-bit quantization to accelerate runtime decoding, but the approach fundamentally remains column-major.
Bit-serial LUT-based GEMM operates by loading weights corresponding to activation combinations and performing table lookups to compute results. This requires careful consideration of the weight organization in memory. Taking activation groups of size gg=4 as an example, Figure 7 illustrates our design for rearranging the one-bit weight matrics to fit the 128-bit ARM NEON registers and efficiently unpacking them at runtime.
Given ARM’s 128-bit register width, a 16×8 one-bit matrix block stored in conventional row-major or column-major order would reside in non-contiguous memory. To maximize memory bandwidth utilization, we reorganize such blocks along the out-channel dimension, ensuring contiguous memory access. Since weights remain static during inference, this rearrangement can be performed offline, introducing no runtime overhead. During execution, weight decoding only requires simple bitwise AND and shift operations, maintaining computational efficiency.
4.5.2 Mirror Storage.
Given the formulation:
$\hat{\mathrm{W}}=\sum_{j=0}^{q-1}s_{j}\cdot b_{j}+z,b_{j}\in\{0,1\}^{n},$ (9)
we apply a simple linear transformation by setting $\hat{s}_{j}=\frac{1}{2}s_{j}$, $\hat{b}_{j}=2b_{j}-1$, $\hat{z}=z+\frac{1}{2}\sum_{j=0}^{q-1}s_{j}$, under this transformation, the quantized weights can be rewritten as:
$\mathrm{\hat{W}}=\sum_{j=0}^{q-1}\hat{s}_{j}\cdot\hat{b}_{j}+\hat{z},\hat{b}_{j}\in\{-1,1\}^{n}.$ (10)
For an input activation combinations of size 4, such as $x_{1},x_{2},x_{3},x_{4}]$, the output of the dot product with the weight has 16 possible outcomes, ranging from $(-x_{1}-x_{2}-x_{3}-x_{4},\ ...,\ x_{1}+x_{2}+x_{3}+x_{4})$. When storing the lookup table, we only need to store half of the possible results, as the remaining half can be obtained by negating the stored values. This table compression method is lossless, fully preserving model inference accuracy while also reducing memory usage by half and accelerating table access.
4.5.3 Table Quantization.
For ARM NEON, activations are typically stored in FP16 precision, and accordingly, each entry in the lookup table also uses FP16 precision. To optimize table storage and access, each FP16 value can be split into two int8 values and interleaved in memory. During table loading, efficient dual-channel load operations such as vld2q_u8vld2q\_u8 can be employed to construct the table, after which the retrieved values are reconstructed back into FP16.
An alternative approach is to quantize the table itself, mapping FP16 table values to int8. While this may introduce some degradation in model accuracy, it eliminates the need for interleaved storage and FP16 reconstruction, significantly improving runtime efficiency.
5 Experiments
Our experiments consist of three parts. First, we evaluate the performance of the proposed hierarchical linear quantization method under both PTQ and efficient finetuning settings. Second, we assess the time and memory efficiency of our quantization approach. Finally, we evaluate our CPU kernel and the end-to-end inference performance of the quantized model on edge hardware.
5.1 Main Results
5.1.1 Models, Datasets and Baseline.
We conduct experiments on several widely used models, including LLaMA2 (Touvron et al., 2023), LLaMA3 (Dubey et al., 2024), Qwen3 (Yang et al., 2025). We use the C4 (Raffel et al., 2020) dataset as the calibration set. For 3-bit quantization, Our results are compared with weight-only methods employing uniform quantization schemes, such as GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024), and OmniQuant (Shao et al., 2023). For 2-bit quantization, we additionally compare our method with several efficient mixed-precision quantization approaches, including PB-LLM (Shang et al., 2023) and DB-LLM (Chen et al., 2024b).
5.1.2 Evaluation.
We evaluate all models using perplexity on the Wikitext2 (Merity et al., 2016) and C4 dataset. Additionally, for LLaMA2-7B, LLaMA2-13B, LLaMA3-8B, we assess their zero-shot capabilities using the LM Harness framework (Gao et al., 2021). The evaluation tasks include ARC Easy (Clark et al., 2018), ARC Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), and PIQA (Bisk et al., 2020).
5.1.3 Configuration.
We evaluate weight-only quantization at two precision levels: INT3 and INT2, while keeping activations in full precision. For HLQ-GPTQ, following previous PTQ settings, we select 128 calibration samples from the C4 dataset. we set the maximum number of search step $T_{max}$=100, with a learning rate of $1\times 10^{-3}$ and and a convergence threshold of $1\times 10^{-6}$. For efficient finetuning, we randomly select 1K calibration samples from the C4 dataset. In the block-wise stage, we set the learning rate to $1\times 10^{-4}$ and train for 2 epochs, while in the end-to-end stage, the learning rate is set to $2\times 10^{-5}$ with 1 training epochs.
Table 1: A comparison of perplexity ($\downarrow$) between weight-only quantization methods on C4 and WikiText-2 datasets, with a context length of 2048. Wbits denotes the bit-width of weights, while BPW represents the average number of bits per weight. Scale and zero-point are assumed to be stored in fp16 format. PB-LLM* denotes the result of PB-LLM with GPTQ and 20% salient weight. ”–” indicates that the framework does not support this model.
Method # W # G BPW LLaMA2-7 LLaMA2-13 LLaMA3-8 Qwen3-8
wikitext2 c4 wikitext2 c4 wikitext2 c4 wikitext2 c4
Baseline 16 - 16 5.47 6.97 4.88 6.47 6.15 8.89 9.72 13.30
GPTQ 2 128 2.25 16.01 33.70 10.33 20.97 109.30 181.82 29.19 35.57
AWQ 2 128 2.25 2.21e5 1.72e5 1.23e5 9.41e4 1.74e6 2.14e6 1.21e5 2.52e6
OmniQuant 2 128 2.25 11.06 15.02 8.26 11.05 18.50 35.73 - -
PB-LLM* - - 2.2 17.19 20.60 12.47 15.32 21.84 57.33 - -
HLQ-GPTQ 2 128 2.37 9.90 14.89 7.02 9.75 17.32 27.14 18.13 24.60
HLQ-Finetuning 2 128 2.37 8.05 11.21 6.75 9.43 10.95 19.08 17.12 22.14
DB-LLM 2 64 - 7.23 9.62 6.19 8.38 13.60 19.20 - -
HLQ-GPTQ 2 64 2.75 8.72 13.27 6.47 9.24 13.22 20.52 17.13 20.82
HLQ-Finetuning 2 64 2.75 7.07 9.12 6.07 8.02 10.14 17.08 15.71 18.24
GPTQ 3 128 3.25 6.29 7.89 5.42 7.00 9.58 11.66 10.76 14.39
AWQ 3 128 3.25 6.24 7.84 5.32 6.94 8.16 11.49 14.90 18.51
OmniQ 3 128 3.25 6.03 7.75 5.28 6.98 8.27 11.66 - -
HLQ-GPTQ 3 128 3.5 5.97 7.66 5.14 6.90 7.41 10.85 10.54 14.15
HLQ-Finetuning 3 128 3.5 5.93 7.55 5.12 6.83 7.26 10.74 9.95 13.54
5.1.4 Perplexity Results.
Table 1 presents the perplexity results on the C4 and WikiText-2 datasets.
For HLQ-GPTQ, our method enhances the performance of GPTQ and surpasses other PTQ approaches with uniform quantization. Under the 3-bit configuration, compared to GPTQ, our approach achieves a perplexity reduction of approximately 0.3 on WikiText-2 for both Llama2-7B and Llama2-13B, and a reduction of 2.17 for Llama3-8B. On the C4 dataset, perplexity is reduced by about 0.15 for the Llama2 models and by 0.8 for Llama3-8B. In the 2-bit setting, our method substantially improves upon GPTQ, particularly for the Llama3-8B model, where perplexity on WikiText-2 is reduced from 109.30 to 17.32, and on C4 from 181.82 to 27.41. Although 2-bit results do not surpass those of DB-LLM, we consider this reasonable. DB-LLM incorporates a complex knowledge distillation process, which demands substantial computational resources and time for finetuning. Moreover, its quantization format is less amenable to practical hardware deployment. In contrast, our method preserves the simplicity of PTQ, requiring minimal computational resources and time to complete model quantization.
For HLQ-Finetuning, our method further improves model performance under low-bit quantization, particularly in the 2-bit setting. For the W2g128 configuration of Llama3-8B, the perplexity on WikiText2 and C4 is further reduced to 10.95 and 19.08, respectively. Moreover, under the W2g64 configuration, our method surpasses DB-LLM, while requiring only 1K calibration samples compared to 20K used by DB-LLM.
5.1.5 Zero-Shot Tasks.
Table 2 presents the experimental results on five zero-shot tasks. Under 3-bit quantization, HLQ-GPTQ and HLQ-finetuning all improve the average accuracy by approximately 0.9% compared to GPTQ on Llama2-7B and by approximately 5% on Llama3-8B. For 2-bit quantization, HLQ-GPTQ achieves an improvement of roughly 9% over GPTQ on Llama3-8B, also outperforming AWQ and OmniQuant, while HLQ-Finetuning achieves an improvement of 15%.
Table 2: Accuracy (%) on five zero-shot tasks. at. PB-LLM* denotes the result of PB-LLM with GPTQ and 20% salient weight. Bold: best result; underlined: second-best. ”–” indicates that the framework does not support this model.
Model Method Bits Group WinoGrande HellaSwag ArcC ArcE PiQA Average($\uparrow$)
LLaMA2-7B Baseline 16 - 69.22 57.16 43.52 76.26 78.07 64.85
GPTQ 3 128 68.59 53.66 40.19 73.74 76.01 62.44
AWQ 3 128 67.40 54.98 41.64 74.07 76.01 62.82
OmniQ 3 128 66.69 54.42 39.85 74.37 76.77 62.42
HLQ-GPTQ 3 128 68.11 55.29 41.75 74.25 77.18 63.32
HLQ-Finetuning 3 128 69.14 55.35 41.92 74.33 76.93 63.52
GPTQ 2 128 55.17 32.59 21.25 40.45 58.32 41.56
AWQ 2 128 50.00 26.52 26.79 26.14 49.64 35.82
OmniQ 2 128 55.88 40.28 23.46 50.13 65.13 46.98
PB-LLM* - - 50.36 30.49 22.01 29.88 55.22 37.60
HLQ-GPTQ 2 128 59.67 38.61 25.26 54.34 67.30 49.04
HLQ-Finetuning 2 128 62.75 46.61 30.72 63.01 72.25 55.07
LLaMA2-13B Baseline 16 - 72.22 60.07 48.29 79.42 79.05 67.81
GPTQ 3 128 70.88 57.83 45.65 77.99 78.56 66.18
AWQ 3 128 71.82 58.58 44.62 77.95 77.75 66.14
OmniQ 3 128 70.01 58.46 46.16 77.86 78.40 66.18
HLQ-GPTQ 3 128 70.95 58.10 45.79 78.76 78.60 66.44
HLQ-Finetuning 3 128 71.94 58.54 45.32 78.70 78.56 66.62
GPTQ 2 128 55.80 41.06 21.93 55.60 67.08 48.29
OmniQ 2 128 57.93 46.23 30.29 63.22 70.13 53.56
PB-LLM* - - 52.33 30.23 23.12 31.27 55.01 38.39
HLQ-GPTQ 2 128 67.43 47.83 35.14 66.92 74.10 58.28
HLQ-Finetuning 2 128 66.38 50.78 36.18 67.72 75.19 59.25
LLaMA3-8B Baseline 16 - 72.61 60.17 50.43 80.09 79.60 68.58
GPTQ 3 128 70.88 55.13 37.80 65.24 73.83 60.58
AWQ 3 128 70.96 55.43 44.20 75.84 77.69 64.82
OmniQ 3 128 - - - - - -
HLQ-GPTQ 3 128 71.11 56.01 42.94 76.81 78.35 65.04
HLQ-Finetuning 3 128 69.25 55.75 45.65 77.31 78.13 65.22
GPTQ 2 128 47.91 27.40 19.37 28.45 54.52 35.53
OminQ 2 128 - - - - - -
HLQ-GPTQ 2 128 59.27 38.96 22.18 40.70 60.66 44.35
HLQ-Finetuning 2 128 56.83 41.42 27.13 58.25 69.12 50.55
Table 3: Comparison of Model Compression Rates between Uniform Quantization and Hierarchical Linear Quantization. Embedding and LM-head layers are excluded from quantization for all models. The symbol ’-’ denotes per-channel quantization. ”Rates($\uparrow$)” denote the compression ratio, and ”Mem($\downarrow$)” indicates the size of the compressed model.
Model Wbit G Uniform HLQ
Rates Mem(GB) Rates Mem(GB)
LLaMA2-7B 2 - 6.26 2.002 6.25 2.005
2 128 5.74 2.185 5.50 2.279
3 - 4.55 2.756 4.54 2.762
3 128 4.27 2.938 4.01 3.127
LLaMA3-8B 2 - 4.17 3.588 4.16 3.592
2 128 3.95 3.785 3.84 3.887
3 - 3.40 4.401 3.39 4.407
3 128 3.25 4.598 3.11 4.801
LLaMA2-13B 2 - 6.78 3.574 6.77 3.579
2 128 6.16 3.934 5.88 4.119
3 - 4.80 5.051 4.79 5.060
3 128 4.48 5.411 4.19 5.780
5.1.6 Compression Rates.
It can be observed that under the same weight precision setting, HLQ incurs a higher average bit-width compared to uniform quantization. For example, at 2-bit, the average bit-width is 2.25 for uniform quantization versus 2.37 for HLQ; at 3-bit, it is 3.25 for uniform quantization versus 3.5 for HLQ. This is because HLQ requires storing a separate scale for each bit plane. Table 3 presents the model compression rates. For LLaMA3-8B, HLQ at 2-bit requires an additional 0.1 GB of storage space compared to uniform quantization—an increase of approximately 5%. However, this storage overhead does not introduce additional computational cost in Bit-Serial LUT-based GEMM, enabling end-to-end inference performance on par with uniform quantization. We will provide a detailed analysis of this aspect in Section 9.
5.2 Ablation Study.
5.2.1 Hyparameter Study for HLQ-GPTQ.
We conducted a hyperparameter search for HLQ-GPTQ, and the results are summarized in Table 4. For both 2-bit and 3-bit quantization, we evaluated three different learning rates: 1e-2, 1e-3, and 1e-4. Our experiments indicate that for 3-bit quantization and 2-bit quantization , a learning rate of 1e-3 is recommended. Furthermore, we observed that the learning rate has a significant impact on the final model performance. With a carefully selected learning rate, our method is able to achieve improved results. The use of regularization techniques is a promising direction for achieving stable performance across varying learning rates, which we leave for future work.
Table 4: Hyparameter study for HLQ-GPTQ, c4 PPL($\downarrow$) on LLaMA3-8B is reported.
Learning rate 1e-2 1e-3 1e-4
3 bit 1.4e3 10.85 11.17
2 bit 27.50 27.14 50.04
5.2.2 Calibration Data Size for Efficient Finetuning.
We investigated the effect of the size of the calibration data on the final performance. As illustrated in Figure 8, our method yields substantial improvements as the calibration set size increases when it is below 1k. However, once the size exceeds 1k, the improvement becomes marginal. Since enlarging the calibration set introduces considerable memory overhead, especially during the block-wise stage, we set the calibration set size to 1k, thereby striking a balance between performance and efficiency.
Figure 8: Model performance of w2g128 Llama3-8b with different calibration size when finetuning.
5.2.3 Impact of Each Components in Efficient Finetuning.
Finetuning consists of two main stages: block-wise quantization and end-to-end tuning. We investigate the impact of each stage on the final model performance. The results are presented in Table 5.
Table 5: Impact of each component in Finetuning on Llama3-8B w2g128 quantization.
Block-Wise E2E-Finetuning c4 PPL($\downarrow$) avg.acc($\uparrow$)
1937.08 30.11
27.47 50.07
583.75 35.42
19.08 50.55
5.3 Efficiency Evaluation.
5.3.1 Time and Memory Efficiency.
Table 6 presents the efficiency of our method in terms of both time and memory. For HLQ-GPTQ, we retain the high computational and memory efficiency of GPTQ in quickly completing model quantization.
In terms of time, for LLaMA3-8B, it requires only 0.7 hour to finish quantization. During the efficient finetuning stage, although parameter training is involved, quantization can still be completed within two hours. A more detailed comparison with other efficient methods is provided in Table 7. Compared with GPTQ, HLQ-GPTQ introduces an additional 0.25 hours of overhead due to parameter search, yet it still maintains high efficiency, achieving a runtime comparable to AWQ. Compared with OmniQuant, which also requires block-wise reconstruction, HLQ-Finetuning has a significant advantage in quantization time, as it iterates only 2 epochs per block and requires merely 27% of the time needed by OmniQuant to complete quantization. In terms of memory, our method achieves very high efficiency since only a small subset of parameters are trained.
Table 6: Quantization time and peak memory usage of HLQ-GPTQ and HLQ-Finetuning. All experiments are conducted on an NVIDIA RTX 4090 GPU.
Model HLQ-GPTQ HLQ-Finetuning
T Mem(3/2 bit) T Mem(3/2 bit)
llama2-7 0.5h 8/7GB 1.5h 10/9GB
llama3-8 0.7h 9/8GB 2h 13/10GB
llama2-13 1.2h 10/9GB 4.5h 16/12GB
Table 7: Comparison of quantization time and peak memory usage on W2g128 Llama2-7B across different efficient quantization methods.
Method Time(h) Memory(GB)
GPTQ 0.20 8
AWQ 0.50 4
HLQ-GPTQ 0.50 8
OmniQ 5.50 10
HLQ-Finetuning 1.50 9
5.3.2 Data Efficiency.
Compared with previous approaches, our method demonstrates substantially higher data efficiency. For HLQ-GPTQ, we follow prior works and use only 128 calibration samples. For HLQ-Finetuning, although our final experiments adopt a calibration set size of 1k, even with only 128 samples our method still outperforms previous uniform quantization methods. In contrast, non-uniform quantization methods such as DB-LLM require up to 20K calibration samples, while our approach achieves competitive performance with only 5% of their requirement—thereby significantly reducing the calibration data demand.
5.4 Hardware Deployment on Edge Devices
5.4.1 Hardware Devices.
We evaluate the inference performance of our framework on edge devices based on the ARM architecture. ARM-based chips typically adopt a highly integrated System-on-Chip (SoC) design and cache-friendly, which are advantageous for table lookup operations. In addition, ARM processors are known for their low power consumption, making them widely used in edge computing scenarios. For evaluation, we select two representative ARM-based chips: Apple’s M2 chip from the Apple Silicon series and the RK3588 chip based on the ARM Cortex series.
5.4.2 Evaluation Setup.
In terms of hardware evaluation, our experiments focus on the following aspects:
  • Numerical precision evaluation: Since our optimization algorithm is executed on the GPU, deploying the optimized model on edge devices may introduce numerical precision errors. Therefore, we evaluate the numerical accuracy of our framework. The experimental results demonstrate that the output of our framework achieves a cosine similarity greater than 99% compared to the GPU output, meeting the requirements for practical application.
  • Overhead of HLQ: we follow T-MAC’s design and adapt it to support HLQ. Experimental results show that the introduction of HLQ incurs no additional overhead to the Bit-serial LUT-based GEMM kernel.
  • End-to-End Speedup Comparison with AWQ: We compare our inference framework with AWQ for end-to-end speedup.
  • Compared with llama.cpp: We further compare our framework with the state-of-the-art edge-side inference framework, llama.cpp, to demonstrate the superiority of our method.
5.4.3 Numerical Precision Evaluation.
As mentioned in TensorRT (NVIDIA, 2023) and the MLPerf (Reddi et al., 2020), the cosine similarity between GPU and CPU outputs needs to be greater than 99% to meet the application standards. For the LLaMA2-7B model with 3-bit weight quantization, we select the output hidden states from the 9th, 18th, and 31st (final) layers on the M2 chip and compare them with the corresponding outputs on the GPU. The hidden states of the GPU denote $\mathbf{a}$, and the hidden states of the M2 denote $\mathbf{b}$. The cosine similarity is computed as
$\text{cos\_sim}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$
Results in Table 8 demonstrate that the cosine similarity between our framework and the GPU simulation exceeds 99.9%.
Table 8: Cosine similarity(%) between simulated quantization on GPU and actual quantization on the M2 chip at the 9th, 18th, and 31st(final) layers of the LLaMA2-7B model with 3-bit weight quantization.
Model Prompt L9 L18 L31
LLaMA2-7B-3Bit 128 99.9994 99.9990 99.9982
256 99.9984 99.9986 99.9970
512 99.9988 99.9978 99.9960
5.4.4 Overhead of HLQ.
We modified the T-MAC operators to support the HLQ format and evaluated their performance at the kernel level. Table 9 presents the latency results of two types of matrix operations in the 2-bit quantized LLaMA2-7B model, evaluated on both the RK3588 and Apple M2 chips. For each operation, the reported latency is the average of 10 runs. The results demonstrate that the introduction of HLQ does not introduce additional overhead to the Bit-serial LUT-based GEMM. The increase in average latency is negligible compared to the inherent fluctuations and overall runtime. Therefore, although the HLQ quantization format occupies slightly more memory than uniform quantization under the same bit-width and group configuration, it does not introduce any additional computational overhead during inference. Moreover, the extra memory required by HLQ is small relative to the overall size of the compressed model, typically remaining below 5%, which is considered acceptable in practice.
Table 9: Comparison of GEMM latency (ms) between uniform quantization and HLQ formats using the T-MAC kernel.
Chip Kernel $11008\times 4096$ $4096\times 32000$
RK3588 T-MAC-Uniform 5.70(± 0.10)5.70\,($\pm\,$0.10) 22.11(± 0.17)22.11\,($\pm\,$0.17)
T-MAC-HLQ 5.72(± 0.11)5.72\,($\pm\,$0.11) 22.23(± 0.16)22.23\,($\pm\,$0.16)
M2 T-MAC-Uniform 1.84(± 0.07)1.84\,($\pm\,$0.07) 4.46(± 0.05)4.46\,($\pm\,$0.05)
T-MAC-HLQ 1.86(± 0.04)1.86\,($\pm\,$0.04) 4.49(± 0.05)4.49\,($\pm\,$0.05)
Table 10: End-to-end performance comparison of ELUTQ and AWQ on LLaMA3-8B running on Apple M2 (batch size = 1, prompt length = 512).
Method W G Mem(GB) Speedup
[…]
Table 11: Comparison of prefill and decode throughput (tokens/s) between ELUTQ and llama.cpp on the Apple M2 chip. All experiments are conducted with 4 threads and a batch size of 1. More results can be found in Appendix B
Model Framework Wbits BPW Input,Output Prefill Decode
llama2-7B llama.cpp Q3_K_M 3.91 128,128 19.87 15.65
256,128 19.75 15.53
Q3_K_S 3.50 128,128 17.60 15.44
256,128 17.43 15.00
Q2_K_S 2.50 128,128 14.82 13.35
256,128 14.54 13.16
ELUTQ W3(g128) 3.50 128,128 20.60 16.63
256,128 21.54 16.21
W2(g128) 2.37 128,128 25.06 20.48
256,128 25.94 20.25
5.4.5 End-to-End Speedup Comparison with AWQ
Here, we compare ELUTQ with the uniform-quantization-based inference framework AWQ in terms of end-to-end speedup on the LLaMA3-8B model. All experiments are conducted on the Apple M2 chip. As shown in Table 10, under the 3-bit setting, AWQ achieves a 2.0× speedup over FP16, while ELUTQ achieves 2.5×. Under the 2-bit setting, AWQ achieves 1.6×, whereas ELUTQ reaches 3.4×. These results indicate that AWQ suffers from significant dequantization overhead at lower bit widths, resulting in reduced speedup at 2 bits, whereas our framework continues to improve efficiency as the bit width decreases.
5.4.6 Compared with llama.cpp.
Directly comparing inference speed with llama.cpp is not entirely fair, since llama.cpp adopts its own quantization format, GGUF, which is incompatible with our quantization scheme. To ensure a fair comparison, we instead use bits per weight (BPW) to measure model size, focusing on how inference speed changes as the model size varies.
Table 11 presents the results on LLaMA2-7B and Qwen2-1.5B. We observe that under lower-bit settings, llama.cpp becomes slower as the model size decreases. For example, Q3_K_S is slower than Q3_K_M and Q2_K_S is slower than Q3_K_S. This phenomenon mainly arises from the substantial decoding overhead associated with low-bit quantization. In contrast, our framework benefits from the Bit-serial LUT-based GEMM paradigm, achieving nearly linear improvements in inference speed as the average bit-width decreases. Specifically, when BPW = 3.5, ELUTQ outperforms llama.cpp by approximately 20% in the prefill stage and 10% in the decode stage. When BPW $\approx$ 2.5, the speedup increases to about 45% and 25%, respectively.
Moreover, the improvement is more pronounced in the prefill stage than in the decode stage. This is because the prefill stage is computation-intensive and primarily involves GEMM operations, while the decode stage is memory-intensive and mainly consists of GEMV operations. In the prefill stage, our method replaces multiplication in matrix multiplication with efficient lookup and sum operations, significantly accelerating computation. In contrast, the decode stage is memory-intensive, where runtime is dominated by memory access. Since the Bit-serial LUT-based GEMM requires additional table lookups, its memory access cost is slightly higher than that of traditional GEMV, leading to smaller speed gains in the decode stage.
6 Limitations and Discussion
Here, we discuss the limitations of our work and outline several promising directions for future research.
  • Weight-only quantization. The current study focuses exclusively on weight only quantization. Extending our framework to weight activation quantization could further enhance inference efficiency and reduce memory consumption. This remains an open and valuable direction for future research.
  • Integration with other quantization frameworks. At present, HLQ has been integrated only with a limited set of efficient quantization methods. Combining HLQ with Quantization-aware Training or LoRA-based adaptation may lead to additional gains in quantization accuracy, which we plan to explore in subsequent work.
  • Optimization dependency and stability. The quantization parameters in HLQ, including scale and zero point, are currently optimized through gradient based search. Consequently, the final performance is somewhat sensitive to the choice of learning rate. Future work could incorporate interpretable regularization strategies or adaptive optimization mechanisms to improve the stability and robustness of the training process.
  • Hardware generalization. HLQ also shows potential for broader deployment across various hardware platforms such as GPUs and NPUs. Extending the method to these architectures represents another promising avenue for future investigation.
7 Conclusion
In this paper, we propose ELUTQ, an efficient quantization framework designed for deploying large language models on edge devices. The framework is carefully aligned with the computation pipeline of Bit-serial LUT-based GEMM and introduces a novel Hierarchical Linear Quantization method that better captures the weight distribution compared to traditional uniform quantization. Moreover, HLQ can be seamlessly integrated with existing quantization techniques, including post-training and fine-tuning, to further enhance model accuracy. Finally, we develop a pure C++ inference framework that enables accurate and efficient on-device inference, facilitating the practical deployment of quantized models in edge scenarios.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 92473208), the Tianjin Science and Technology Planning Program (Grant No. 24ZYCGYS00680), and in part by the Tianjin Key Laboratory of Optical-electronic Sensor and Sensor Network Technology.
Appendix A Alternating Optimization
Alternating optimization is a heuristic algorithm. At initialization, we set $z=\min(\mathrm{W})$, and the initial value of s is set to the scale factor used in uniform quantization. Specifically, let $\Delta=\frac{\max(\mathrm{W})-\min(\mathrm{W})}{2^{q}-1}$, For a $q$-bit quantization, the initial value of s is set to $s=[\Delta,2\Delta,...2^{q-1}\Delta]$. After initialization, we solve the minimization problem in formulation 7 through an alternating optimization procedure between Bit-Pattern Selection and Linear Reconstruction.
Bit-Pattern Selection. At $t$-th step, given the current scale parameters $s^{(t-1)}$ and zero-points $z^{(t-1)}$, we determine the optimal bit pattern $B^{(t)}$ that minimizes the reconstruction error under the fixed quantization parameters:
$B^{(t)}=\arg\min_{B}\|\mathrm{W}-\mathrm{Dequant}(B,s^{(t-1)},z^{(t-1)})\|^{2},$ (11)
where $\mathrm{W}$ denotes original pretrained weight, $\mathrm{Dequant}$ denotes dequantization method. This process is essentially equivalent to the one described in Equation 5, i.e., for each $w$, selecting the optimal bit combination from the candidate codebook.
Linear Reconstruction. Once the optimal bit pattern $B_{t}$ is obtained, we fix $B_{t}$ and update the continuous quantization parameters by solving a least-squares regression problem:
$(s^{(t)},z^{(t)})=\arg\min_{s,z}\|\mathrm{W}-(B^{(t)}s+\mathbf{1}\cdot z)\|^{2}.$ (12)
Here, adding a constant column of ones allows the zero-point to be incorporated directly into the linear system. This formulation enables joint optimization of scales and zero-point given a fixed discrete structure and reduces to a standard linear least-squares problem with a closed-form solution. Therefore, the optimization objective in Equation 12 can be equivalently expressed as follows:
$(s^{(t)},z^{(t)})=\mathrm{LSE}{(\mathrm{W},B^{(t)})},$ (13)
where $\mathrm{LSE}$ denotes the standard least squares estimation.
We summarize alternating optimization in Algorithm 2.
Algorithm 2 Alternating Optimization for Hierarchical Linear Quantization
0: Weight matrix $\mathrm{W}\in R^{n\times k}$, bit width q, Group size g, Max iterations $T_{max}$
0: scales $s\in R^{n\times\frac{k}{g}\times q}$, zero points $z\in R^{n\times\frac{k}{g}}$
1: Reshape W into groups $\mathrm{\widetilde{W}}\in R^{n\times\frac{k}{g}\times g}$
2: Calculate uniform quantization scale : $\Delta\leftarrow\frac{\max(\mathrm{\widetilde{W}})-\min(\mathrm{\widetilde{W}})}{2^{q}-1}$
[…]
4: Generate binary combinations $C$
5:for $t\leftarrow 1$ to $T_{max}$ do
6:  $\mathrm{\hat{W}^{(t)}}\leftarrow\mathrm{\widetilde{W}}-z^{(t-1)}$
7:  $V^{(t)}\leftarrow s^{(t-1)}\times C^{T}$ # Matrix product
8:  $k^{*}\leftarrow\arg\min_{k}\|\mathrm{\hat{W}^{(t)}}-V^{(t)}_{k}\|$
9:  $B^{(t)}\leftarrow C_{k^{*}}$ # Choose best binary combination
10:  $s^{(t)},\ z^{(t)}\leftarrow LSE(\mathrm{\widetilde{W}},B^{(t)})$ # Least Squares Estimation
11:end for
We conduct an experiment where HLQ-GPTQ is implemented using both alternating optimization and gradient-based search methods. As shown in Table 12, the gradient-based optimization generally outperforms the alternating approach in most cases and exhibits greater stability. Therefore, considering algorithmic robustness, we recommend adopting the gradient-based method in practical applications. Nevertheless, the alternating optimization approach can complete model quantization within a very short time (Table 11) and achieves better performance than GPTQ under low-bit configurations.
Table 12: Performance comparison of HLQ-GPTQ using alternating optimization and gradient-based optimization across different models and bit configurations. perplexity($\downarrow$) is reported.
Method # W # G BPW LLaMA2-7 LLaMA2-13 LLaMA3-8 Qwen3-8
wikitext2 c4 wikitext2 c4 wikitext2 c4 wikitext2 c4
Baseline 16 - 16 5.47 6.97 4.88 6.47 6.15 8.89 9.72 13.30
Alternating Optimization 2 128 2.37 15.72 29.81 9.83 16.45 14.32 25.54 18.43 25.02
Gradient-based Method 2 128 2.37 9.90 14.89 7.02 9.75 17.32 27.14 18.13 24.60
Alternating Optimization 3 128 3.5 6.52 8.43 5.73 7.26 7.35 10.88 10.73 14.62
Gradient-based Method 3 128 3.5 5.97 7.66 5.14 6.90 7.41 10.85 10.54 14.15
Table 13: Runtime (h) comparison of GPTQ, HLQ-GPTQ (Alternate), and HLQ-GPTQ (Gradient).
Model GPTQ HLQ-GPTQ(Alternate) HLQ-GPTQ(Gradient)
LLaMA2-7B 0.20 0.25 0.50
LLaMA2-13B 0.30 0.40 1.00
LLaMA3-8B 0.20 0.30 0.75
Appendix B More results on Edge Devices
Table 14: Comparison of prefill and decode throughput (tokens/s) between ELUTQ and llama.cpp on the Apple M2 chip and RK3588. All experiments are conducted with 4 threads and a batch size of 1.
Model Framework Wbits BPW Input,Output Apple M2 RK3588
Prefill Decode Prefill Decode
Qwen2-1.5B llama.cpp Q3_K_M 3.91 128,128 103.11 57.69 21.53 15.36
256,128 101.37 56.11 21.28 15.13
Q3_K_S 3.50 128,128 86.67 54.64 19.79 14.57
256,128 85.54 53.49 19.58 14.37
Q2_K_S 2.50 128,128 62.73 48.55 16.12 12.95
256,128 60.17 47.87 15.98 12.78
ELUTQ W3(g128) 3.50 128,128 88.72 55.03 27.77 14.84
256,128 93.06 54.90 27.16 14.54
W2(g128) 2.37 128,128 120.65 65.19 33.61 17.42
256,128 124.36 64.81 35.39 17.15